OpenStudy (anonymous):

Write a simplified polynomial expression to represent the area of the rectangle below. 2x-4 x+5

OpenStudy (anonymous):

Can you combine any 'like-terms' ?

OpenStudy (anonymous):

r u asking me to answer that because you know the answer and you want me to answer it or are u asking because you dont know??

OpenStudy (anonymous):

I know you can, I'm asking if you know how to combine them.

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

1x+2x =3x, so 2x-4x = ?

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

but how did you get 1x

OpenStudy (anonymous):

i'm just showing you the form In general: 1x+2x=3x, 5x-3x=2x In your specific problem 2x-4x = ? (it's not 6x)

OpenStudy (anonymous):

idk the answer but i luv ur pic puertoricangirl

OpenStudy (anonymous):

@ nicole lol thank

OpenStudy (anonymous):

you

OpenStudy (anonymous):

if its not 6x then it has to be 8x

OpenStudy (anonymous):

2x-4x+5 can be re-written as x(2-4)+5, can you solve this?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

what is 2-4?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

like in my example x(2-4)+5, would yield x(-2)+5

OpenStudy (anonymous):

Sorry, so your equation 2x-4x would = -2x

OpenStudy (anonymous):

So once you make that combination, what are you left with in your equation?

OpenStudy (anonymous):

You started with 2x-4x+5, we just worked the 2x-4x = -2x, so you're left with -2x+5

OpenStudy (anonymous):

ok then what...

OpenStudy (anonymous):

that's it, -2x+5 is all you do. It asked for a polynomial expression, which means it will have an x in it.

OpenStudy (anonymous):

oooooooooooooooooohhhhhhhhhhhhhhhh!!!!!!!!!!!!!!!!!!